# Multi-neutron correlations in light nuclei via ab-initio lattice simulations

### v1

@@ Abstract
The quest to understand multi-neutron systems has a long history, and recent experimental efforts aim to probe candidate four-neutron configurations in neutron-rich light nuclei such as 8He and 7H via quasi-free knockout reactions. However, the ground-state energies of the hydrogen isotopes 6H and 7H are not yet well constrained, with substantial discrepancies across experimental analyses and theoretical predictions. Using ab initio nuclear lattice effective field theory with an ensemble of 282 chiral two- and three-nucleon forces, we perform an uncertainty-quantified analysis of the ground-state energies of 6H and 7H. The marginal posteriors suggest single-neutron separation energy MeV, disfavoring sequential decay and pointing towards direct emission. Intrinsic densities indicate triton- and -like clusters in 7H and 8He, respectively. By computing two-body and reduced four-body correlation functions, we find that the valence neutrons in the surface region of these systems form compact dineutrons that predominantly organize into symmetric dineutron-dineutron configurations, with only a small but non-negligible fraction assembling into more compact tetraneutron-like substructures. In 7H, these components account for roughly 95% and 5% of the sampled four-neutron configurations, respectively, and 8He exhibits a similar hierarchy. For these configurations, we also extract the corresponding spatial and angular correlation patterns among the nucleons. These results provide nuclear-structure insights into the debate surrounding four-neutron clusters and complement ongoing experimental searches for tetraneutron signatures in light nuclei.
@@
Motivated by quasi-free knockout experiments [56, 57, 58], which have revealed evidence for preformed clusters in neutron-rich nuclei [27] and have suggested that quasi-free knockout can access four-neutron configurations that closely resemble the intrinsic subsystem of the projectile nucleus [4], we focus on neutron-rich 7H and 8He. These cluster-core systems with four valence neutrons provide unique laboratories for studying emergent correlations, and our analysis of their internal correlations offers complementary nuclear-structure insights into the long-standing experimental and theoretical controversy surrounding and systems (see Ref. [1] and references therein and see Refs. [59, 4, 61, 60, 62] for recent developments). Here, we present an ab initio investigation of the neutron-rich nuclei 7H and 8He using nuclear lattice effective field theory (NLEFT) [63], with full -body correlations. We identify compact surface dineutrons that organize into two representative four-neutron configurations and quantify the geometry and probability of these configurations from two- and four-neutron correlation functions.
@@
Methods: NLEFT provides a powerful framework to capture the full set of many-body correlations through AFMC simulations, but its application to high-fidelity chiral interactions has been hindered by severe sign problems [63, 64, 65]. With recent advancements via the wavefunction matching method, chiral forces up-to-and-including N3LO can now be treated largely free from sign problems [66]. In this work, we employ a set of N3LO chiral interactions that include both nucleon–nucleon (NN) and three-nucleon (3N) forces. These interactions are obtained through history matching [67, 68, 69, 66] that systematically incorporates uncertainties from NN scattering data, the truncation of the chiral hierarchy, and the binding energies of selected nuclei. The main results are derived from calculations using the best-fit interaction, while the corresponding uncertainties and correlations among observables are quantified using Bayesian posterior predictive distributions (PPD) constructed from an ensemble of 282 non-implausible chiral interactions, as detailed in [70].
@@
To probe observables associated with nuclear density, particularly those related to intrinsic structure and correlation patterns, we employ the pinhole method [71]. This approach enables sampling of the spatial coordinates, spins, and isospins of all nucleons according to the quantum amplitudes of the -body density within the auxiliary-field framework. With these pinhole configurations, a wide range of observables of physical interest can be evaluated, as described in [70]. In the following, expectation values are denoted by . In order to reveal the intrinsic structure of the nucleus, an appropriate alignment scheme for the intrinsic density is often employed, as in Refs. [72, 71, 73, 74, 75]. Here, we identify candidate clusters (such as an particle or a triton) by selecting the most probable grouping of nucleons that matches the cluster configuration within each sampled pinhole configuration. Once the clusters are identified, we realign the density profile with respect to the cluster positions, thereby making the intrinsic spatial structure of the nucleus more apparent. Further technical details are provided in [70].
@@
In a four-body subsystem, correlations span many degrees of freedom, including individual coordinates and all pairwise opening angles, making it difficult to extract clear structural information directly. Building on the two-body correlations and symmetry considerations, we construct reduced projections that capture the essential features of the underlying multi-nucleon correlations. In particular, we define a reduced four-body correlation function
where label the nucleons, the sum runs over pairs of disjoint nucleon pairs and , and the angles and for each pair are defined in the same way as in the two-body case. The angle is the opening angle between the c.m. vectors of the two pairs, while denotes the torsion angle between the relative position vectors of nucleons in each pair. This definition projects the four-body density onto configurations in which two disjoint nucleon pairs exhibit a specified angular geometry, as schematically illustrated in [70].
In practice, we perform AFMC simulations on a periodic cubic lattice with side length of and spatial lattice spacing of fm, corresponding to a momentum cutoff of MeV. The temporal lattice spacing is set to MeV-1 for the Euclidean-time evolution. The initial states are chosen as the product of the single-particle harmonic oscillator (HO) wavefunctions. For the hydrogen isotopes, the HO frequency of the valence neutrons is reduced to generate more diffuse wavefunctions and to accelerate convergence. Calculated energies are extrapolated to the infinite Euclidean-time limit using the formalism outlined in [70]. Finite-volume effects have been explicitly examined, and their impact is found to be well within the statistical uncertainties [70]. In the pinhole calculation, the point-nucleon coordinates are convolved with a Gaussian [71] of width fm to account for the finite size of the nucleons [85]. Each pinhole configuration undergoes 3000 Gaussian smearing iterations to enhance the smoothness and reduce lattice artifacts in the density profiles. Unless stated otherwise, density-related observables are evaluated at a Euclidean projection time of MeV-1, ensuring a balance between convergence and statistical precision, following Refs. [71, 74, 75, 86].
@@ Results:
The calculated energies of 6,7H and 8He are shown in Fig. 1 and compared with experimental data. For 8He, multichannel calculations projected onto the irrep of the cubic group [87, 88] yield energies of MeV and MeV for the and states, respectively, consistent with the experimental values MeV [89] and MeV [90]. In addition, the lattice simulations reveal a dineutron correlation in 8He that emerges and persists at sufficiently long Euclidean times, indicative of configuration mixing beyond the initial pure filling. A detailed discussion of this pattern is given below. This aligns with reaction [91, 92, 93] and structure [94, 53, 54] studies that indicate configuration mixing and deformation in 8He, and with the newly observed cluster-like excitation, for which strong – components have been discussed [90]. Extending to hydrogen isotopes, we find a ground state (g.s.) in 6H, indicating occupancy and a breaking of the subshell closure, which is conducive to dineutron correlations [36]. Experimentally, the reported 6,7H g.s. energies and widths vary across reactions (see Table S2 in [70]), and two advanced theoretical approaches predict 6,7H g.s. energies that differ by about 7 MeV [95, 96], underscoring the sensitivity to the interaction and many-body treatment. Incorporating full -body correlations, we perform a Bayesian analysis over 282 non-implausible chiral interactions to quantify uncertainties. The resulting g.s. energy medians and credible intervals for 6,7H are MeV and MeV, respectively, within the experimental ranges. A quantitative treatment of resonance widths requires a dedicated analysis and is beyond the present scope. The marginal posterior for the one-neutron separation energy of 7H suggests a positive at the 1 level, MeV, which kinematically disfavors sequential decay via 6H and thereby makes direct multi-neutron emission channels comparatively more relevant. These findings, together with experimental and theoretical indications [4, 61] of dineutron clustering in dynamics, underscore the emergence of enhanced multi-neutron correlations in 7H and 8He, from dineutron to possible tetraneutron-like four-body configurations, and set the stage for the correlation analysis that follows.
@@
Fig. 2 presents the intrinsic nucleon densities and two-neutron () correlations in 7H, taken as a representative case of multi-neutron dynamics near the dripline (see [70] for benchmarks in 6He and complementary analysis of 8He). The intrinsic frame is defined with the triton c.m. located on the negative axis, so that the valence neutrons in this convention predominantly appear in the upper (positive-) half of the panels. The nucleon and proton density maps on the and slices reveal a compact triton cluster surrounded by four diffuse valence neutrons. This clustered geometry is consistent with a core–plus–valence-neutron picture as commonly employed in cluster and many-body approaches. The interpretation is further supported by Fig. 2(b), which displays pair correlations for all six neutrons relative to the triton c.m. and reveals two distinct components. The inner component is concentrated at short distances with only mild angular modulation, analogous to the correlations in 3H [70].
@@
The outer component emerges at larger radii, where pronounced enhancements are observed at both small and large opening angles between the two neutrons. These two distinct components indicate that our core–valence partition has only a minor impact on the subsequent correlation analysis and, in particular, that it allows us to analyze the correlations among the valence neutrons. Motivated by this clear separation, we extract the two-body correlations among the valence neutrons in the 7H c.m. frame, Fig. 2(c), as commonly used in experimental analyses, and resolve more detailed features of the valence-neutron angular correlations. The resulting bimodal angular pattern closely resembles that obtained in Fig. 2(b), indicating the robustness of the dineutron correlations, and shows pronounced maxima around and at fm. At small , the enhancement indicates the presence of compact dineutron-like pairs in the surface region. We further extract the typical size for pairs with small , as detailed in Refs. [70]. The comparable strength at large contrasts with systems such as 6He, where only a relatively weak large- component has been reported [34, 36, 97]. For 7H, each pinhole configuration contains several neutron pairs, so the two-body correlation function represents a statistical distribution over all pairs and thus provides a probe sensitive to the underlying multi-neutron geometry. The enhanced large- component indicates that, besides compact surface dineutrons, neutron pairs also populate more extended, back-to-back geometries. This pattern is qualitatively consistent with configurations in which the four valence neutrons form two dineutron-like clusters located on opposite sides of the core, in line with a more symmetric – halo geometry inferred from experimental analyses of 8He [98] and supported by microscopic calculations of 8He and 7H that exhibit – clustering [47, 48, 49, 50, 51, 52, 53, 54]. Such bimodal angular correlations are strongly suggestive of – clustering and may represent a microscopic mechanism underlying enhanced correlations in 7H and 8He.
@@
The structure of configurations has been explored analytically using simple schematic models [99, 100] and in intuitive cluster descriptions that treat the four neutrons as two correlated dineutron clusters [47, 48, 49, 50, 51, 52, 53, 54]. However, a systematic characterization of four-neutron geometries at the nucleonic level remains limited, and an explicit four-body treatment is required to access detailed spatial and angular information. In practice, we evaluate Eq. (2) using pinhole samples with two spin-up and two spin-down valence neutrons, which constitute the bulk of the pinhole ensemble. Motivated by the dineutron correlations identified in 7H and 8He, we characterize representative four-neutron configurations in terms of two opposite-spin pair coordinates, which are sensitive to dineutron-like structures. Fig. 3 summarizes the resulting four-neutron geometries. We first present in Fig. 3(a) the marginal distribution of inter-pair opening and torsion angles (, ), obtained by integrating the four-body correlation function defined in Eq. (2) over the intra-pair angles (, ) and (, ). Guided by the correlations of 7H discussed above, we identify dineutron-like configurations by requiring pairs to satisfy and when constructing the marginal distribution of the inter-pair opening and torsion angles, with . The deviation of from quantifies the degree of radial asymmetry within each pair. The robustness of these angular distributions with respect to variations in the cuts is demonstrated in Ref. [70], where the normalized marginal four-body correlation distribution changes only weakly and remains consistent within uncertainties. The marginal probability density is strongly enhanced for opening angles in the range –, and forms a broad maximum centered around . This distribution is most naturally associated with configurations containing two dineutron-like pairs on opposite sides of the system c.m., with finite torsion indicating non-planarity. While Ref. [100] suggested possible coplanar “great-circle” configurations in 8He based on a pure shell-model occupation, our results for 7H show little strength near , indicating that planar arrangements are disfavored when many-body correlations are included explicitly. These features, together with the enhanced probability at finite torsion angles, suggest that the inter- geometry itself is modified by many-body correlations. Consequently, the four-neutron subsystem favors an extended, twisted configuration that is approximately spatially symmetric about the system c.m. but non-planar. In addition to this dominant extended configuration, tetraneutron-like geometries corresponding to the region of small opening angles, where the four neutrons reside on the same side of the system c.m., occur with lower probability. As shown in Fig. 3(b), this component accounts for about 5% of the sampled configurations shown in panel (b) and is concentrated at in the range –. Its associated torsion angle also lies predominantly in the range –, indicating a strongly non-coplanar four-neutron geometry (large ) compared to the symmetric – configuration. For comparison, the distribution in 8He is shifted to smaller angles and more strongly suppressed at large , indicating a less prominent symmetric - geometry than in 7H under the same selection. This may reflect core-induced medium effects on the four-neutron configuration, in that the different proton background and core structure modify the environment experienced by the four neutrons.
@@
To further elucidate these two geometries, Figs. 3(c–f) display the internal angular and radial correlations for configurations corresponding to the extended, approximately symmetric – geometry at and the compact tetraneutron-like geometry at (see [70] for additional details). For the extended – geometry, Figs. 3(d,f), each dineutron pair retains a small internal opening angle , with its distribution peaked near , and the two pairs exhibit nearly symmetric distributions of the distances (defined as the distances from each pair c.m. to the system c.m.) that extend to about 4–5 fm. In contrast, the tetraneutron-like geometry, Figs. 3(c,e), shows enhanced probability density near characteristic radial-distance pairs fm and fm from the system c.m., indicating two radially asymmetric pairs located on the same side of the system c.m. Their internal opening-angle distributions are shifted to larger values, with a pronounced maximum near and lying predominantly in the range -. These configurations are also accompanied by large torsion angles with between and . Together, these features lead to a non-coplanar, interlaced geometry. This geometry represents a compact, strongly correlated four-neutron configuration, in which the distinction between the two clusters becomes less pronounced, suggesting enhanced four-body correlations among the four neutrons. It may be viewed as a structural precursor to a possible tetraneutron configuration. Overall, the contrast between these configurations highlights two representative geometries of the correlated four-neutron subsystem in 7H: a spatially symmetric, extended – geometry with large inter-pair separation and finite torsion, which accounts for roughly 95% of the total marginal distribution, and a compact, twisted tetraneutron-like geometry governed by pronounced four-body correlations, which contributes the remaining 5%. The strong dominance of the extended – geometry in the neutron-rich light nuclei studied here implies that experimental searches for tetraneutron-like configurations are likely to receive sizable contributions from symmetric – modes, posing a serious challenge to isolating a genuine tetraneutron signal.
@@
Summary: We perform ab initio NLEFT calculations of the neutron-rich nuclei 6,7H and 8He, including full -body correlations and a Bayesian analysis over an ensemble of chiral interactions. We predict the ground-state energies of the extreme hydrogen isotopes 6,7H and obtain the energies of the benchmark nucleus 8He in agreement with available data, while simultaneously uncovering their internal multi-neutron structure. Intrinsic densities and two-neutron correlation functions reveal a compact triton (alpha) cluster in 7H (8He) surrounded by four valence neutrons that form compact dineutron pairs. Building on this, we construct a reduced four-body correlation function to resolve the geometry of the four-neutron subsystem. We extract the corresponding spatial and angular correlation patterns between each nucleon in two representative geometries. We find that the 7H ground state is dominated (about 95%) by a spatially symmetric, extended – geometry with large inter-pair separation and finite torsion, while a compact, twisted tetraneutron-like configuration with pronounced four-body correlations accounts for the remaining 5%. These results provide a microscopic picture of emergent multi-neutron correlations in 7H and 8He and imply that experimental searches for tetraneutron-like configurations are likely to be contaminated by sizable backgrounds from symmetric – configurations.

### v2

# Multi-neutron correlations in light nuclei via ab initio lattice simulations

@@ Abstract
The quest to understand multi-neutron systems has a long history, and recent experimental efforts aim to probe candidate four-neutron configurations in neutron-rich light nuclei such as 8He and 7H via quasi-free knockout reactions. However, the ground-state energies of the hydrogen isotopes 6H and 7H are not yet well constrained, with substantial discrepancies across experimental analyses and theoretical predictions. Using ab initio nuclear lattice effective field theory with an ensemble of 282 chiral two- and three-nucleon forces, we perform a Bayesian uncertainty-quantified analysis of the ground-state energies of 6H and 7H. The marginal posteriors suggest single-neutron separation energy MeV, which kinematically disfavors sequential decay via and thereby makes multi-neutron emission channels comparatively more relevant. Intrinsic densities indicate triton- and -like clusters in 7H and 8He, respectively. By computing two-body and reduced four-body correlation functions, we find that the valence neutrons in the surface region of these systems form compact dineutrons that predominantly organize into approximately symmetric dineutron-dineutron configurations, with only a small but non-negligible fraction assembling into more compact tetraneutron-like substructures. In 7H, these components account for roughly 95% and 5% of the sampled four-neutron configurations, respectively, and 8He exhibits a similar hierarchy. For these configurations, we also extract the corresponding spatial and angular correlation patterns among the nucleons. These results provide nuclear-structure insights into the debate surrounding four-neutron clusters and complement ongoing experimental searches for tetraneutron signatures in light nuclei.
@@
Motivated by quasi-free knockout experiments [56, 57, 58], which have revealed evidence for preformed clusters in neutron-rich nuclei [27] and have suggested that quasi-free knockout can access four-neutron configurations that closely resemble the intrinsic subsystem of the projectile nucleus [4], we focus on neutron-rich 7H and 8He. These cluster-core systems with four valence neutrons provide unique laboratories for studying emergent correlations, and our analysis of their internal correlations offers complementary nuclear-structure insights into the long-standing experimental and theoretical controversy surrounding and systems (see Ref. [1] and references therein and see Refs. [59, 4, 61, 60, 62] for recent developments). Here, we present an ab initio investigation of the neutron-rich nuclei 7H and 8He using nuclear lattice effective field theory (NLEFT) [63], with full -body correlations. We identify compact surface dineutrons that organize into two representative four-neutron configurations and quantify the geometry and probability of these configurations from two- and four-body correlation functions.
@@
Methods: NLEFT provides a powerful framework to capture the full set of many-body correlations through AFMC simulations, but its application to high-fidelity chiral interactions has been hindered by severe sign problems [63, 64, 65]. With recent advancements via the wavefunction matching method, chiral forces up-to-and-including N3LO can now be treated largely free from sign problems [66]. In this work, we employ a set of N3LO chiral interactions that include both nucleon–nucleon (NN) and three-nucleon (3N) forces. These interactions are obtained through history matching [67, 68, 69, 66] that systematically incorporates uncertainties from NN scattering data, the truncation of the chiral hierarchy, and the binding energies of selected nuclei. The main results are derived from calculations using the best-fit interaction, while the corresponding uncertainties and correlations among observables are quantified using Bayesian posterior predictive distributions (PPD) constructed from an ensemble of 282 non-implausible chiral interactions, as detailed in Supplemental Material (SM) [70].
@@
To probe observables associated with nuclear density, particularly those related to intrinsic structure and correlation patterns, we employ the pinhole method [71]. This approach enables sampling of the spatial coordinates, spins, and isospins of all nucleons according to the quantum amplitudes of the -body density within the auxiliary-field framework. With these pinhole configurations, a wide range of observables of physical interest can be evaluated. Details of the sampling scheme and observable evaluation are described in SM [70]. In the following, expectation values are denoted by . In order to reveal the intrinsic structure of the nucleus, an appropriate alignment scheme for the intrinsic density is often employed, as in Refs. [72, 71, 73, 74, 75]. Here, we identify candidate clusters (such as an particle or a triton) by selecting the most probable grouping of nucleons that is consistent with the cluster spin-isospin quantum numbers and spatial compactness within each sampled pinhole configuration. Once the clusters are identified, we realign the density profile with respect to the cluster positions, thereby making the intrinsic spatial structure of the nucleus more apparent. Further technical details are provided in SM [70].
@@
In a four-body subsystem, correlations span many degrees of freedom, including individual coordinates and all pairwise opening angles, making it difficult to extract clear structural information directly. Building on the two-body correlations and symmetry considerations, we construct reduced projections that capture the essential features of the underlying multi-nucleon correlations. In particular, we define a reduced four-body correlation function,
where label the nucleons, the sum runs over pairs of disjoint nucleon pairs and , and the angles and for each pair are defined in the same way as in the two-body case. The angle is the opening angle between the c.m. vectors of the two pairs, while denotes the torsion angle between the relative position vectors of nucleons in each pair. This definition projects the four-body density onto configurations in which two disjoint nucleon pairs exhibit a specified angular geometry, as schematically illustrated in Fig. S1 of SM [70].
In practice, we perform AFMC simulations on a periodic cubic lattice with side length of and spatial lattice spacing of fm, corresponding to a momentum cutoff of MeV. The temporal lattice spacing is set to MeV-1 for the Euclidean-time evolution. The initial states are chosen as the product of the single-particle harmonic oscillator (HO) wavefunctions. For the hydrogen isotopes, the HO frequency of the valence neutrons is reduced to generate more diffuse wavefunctions and to accelerate convergence. Calculated energies are extrapolated to the infinite Euclidean-time limit using the formalism outlined in SM [70]. Finite-volume effects were explicitly assessed at the present statistical precision and are found to be subdominant compared with statistical uncertainties [70]. We therefore do not perform a volume extrapolation. In the pinhole calculation, the point-nucleon coordinates are convolved with a Gaussian [71] of width fm to account for the finite size of the nucleons [85]. Each pinhole configuration undergoes 3000 Gaussian smearing iterations to enhance the smoothness and reduce lattice artifacts in the density profiles. Unless stated otherwise, density-related observables are evaluated at a Euclidean projection time of MeV-1, ensuring a balance between convergence and statistical precision, following Refs. [71, 74, 75, 86].
@@ Results:
The calculated energies of 6,7H and 8He are shown in Fig. 1 and compared with experimental data. For 8He, multichannel calculations projected onto the irrep of the cubic group [87, 88] yield energies of MeV and MeV for the and states, respectively, consistent with the experimental values MeV [89] and MeV [90]. In addition, the lattice simulations reveal a dineutron correlation in 8He that emerges and stabilizes at sufficiently long Euclidean times, suggesting configuration mixing beyond the initial pure filling. A detailed discussion of this pattern is given below. This aligns with reaction [91, 92, 93] and structure [94, 53, 54] studies that indicate configuration mixing and deformation in 8He, and with the newly observed cluster-like excitation, for which strong – components have been discussed [90]. Extending to hydrogen isotopes, we find a ground state (g.s.) in 6H, suggesting occupancy and a breaking of the subshell closure, which is conducive to dineutron correlations [36]. Experimentally, the reported 6,7H g.s. energies and widths vary across reactions (see Table S2 in SM [70]), and two advanced theoretical approaches predict 6,7H g.s. energies that differ by about 7 MeV [95, 96], underscoring the sensitivity to the interaction and many-body treatment. Incorporating full -body correlations, we perform a Bayesian analysis over 282 non-implausible chiral interactions to quantify uncertainties. The resulting g.s. energy medians and credible intervals for 6,7H are MeV and MeV, respectively, within the experimental ranges. A quantitative treatment of resonance widths requires a dedicated analysis and is beyond the present scope. The marginal posterior for the one-neutron separation energy of 7H suggests a positive at the 1 level, MeV, which kinematically disfavors sequential decay via 6H and thus makes multi-neutron emission channels comparatively more relevant. These findings, together with experimental and theoretical indications [4, 61] of dineutron clustering in dynamics, underscore the emergence of enhanced multi-neutron correlations in 7H and 8He, from dineutron to possible tetraneutron-like four-body configurations, and set the stage for the correlation analysis that follows.
@@
Fig. 2 presents the intrinsic nucleon densities and two-neutron () correlations in 7H, taken as a representative case of multi-neutron dynamics near the dripline (see SM [70], Figs. S2 for benchmarks in 6He and Figs. S3 for a complementary analysis of 8He). The intrinsic frame is defined with the triton c.m. located on the negative axis, so that the valence neutrons in this convention predominantly appear in the upper (positive-) half of the panels. The nucleon and proton density maps on the and slices reveal a compact triton cluster surrounded by four diffuse valence neutrons. This clustered geometry is consistent with a core–plus–valence-neutron picture as commonly employed in cluster and many-body approaches. The interpretation is further supported by Fig. 2(b), which displays pair correlations for all six neutrons relative to the triton c.m. and reveals two distinct components. The inner component is concentrated at short distances with only mild angular modulation, analogous to the correlations in 3H [70].
@@
The outer component emerges at larger radii, where pronounced enhancements are observed at both small and large opening angles between the two neutrons. These two distinct components indicate that our core–valence partition has only a limited impact on the subsequent correlation analysis and, in particular, that it allows us to analyze the correlations among the valence neutrons. Motivated by this apparent separation, we extract the two-body correlations among the valence neutrons in the 7H c.m. frame, Fig. 2(c), as commonly used in experimental analyses, and resolve more detailed features of the valence-neutron angular correlations. The resulting bimodal angular pattern closely resembles that obtained in Fig. 2(b), indicating the robustness of the dineutron correlations, and shows pronounced maxima around and at fm. At small , the enhancement indicates the presence of compact dineutron-like pairs in the surface region. We further extract the typical size for pairs with small , as detailed in SM [70] (Figs. S3). The comparable strength at large contrasts with systems such as 6He, where only a relatively weak large- component has been reported [34, 36, 97]. For 7H, each pinhole configuration contains several neutron pairs, so the two-body correlation function represents a statistical distribution over all pairs and thus provides a probe sensitive to the underlying multi-neutron geometry. The enhanced large- component indicates that, besides compact surface dineutrons, neutron pairs also populate more extended, back-to-back geometries. This pattern is qualitatively consistent with configurations in which the four valence neutrons form two dineutron-like clusters located on opposite sides of the core, in line with a more symmetric – geometry inferred from experimental analyses of 8He [98] and supported by microscopic calculations of 8He and 7H that exhibit – clustering [47, 48, 49, 50, 51, 52, 53, 54]. Such bimodal angular correlations are suggestive of – clustering and may represent a microscopic mechanism underlying enhanced correlations in 7H and 8He.
@@
The structure of configurations has been explored analytically in schematic models [99, 100] and also within cluster descriptions that treat the four neutrons as two correlated dineutron clusters [47, 48, 49, 50, 51, 52, 53, 54]. A systematic characterization of four-neutron geometries at the nucleonic level has been limited, and an explicit four-body treatment is required to access detailed spatial and angular information. In practice, we evaluate Eq. (2) using pinhole samples with two spin-up and two spin-down valence neutrons, which constitute the bulk of the pinhole ensemble. Motivated by dineutron correlations seen in 7H and 8He, we characterize the four-neutron geometry using two opposite-spin pair coordinates sensitive to dineutron-like structure. Fig. 3(a) shows the marginal distribution of the inter-pair opening and torsion angles obtained from Eq. (2) after integrating over the intra-pair angles and . Guided by the correlations of 7H discussed above, we identify dineutron-like configurations by requiring and , where quantifies the deviation from radial symmetry within each pair. The resulting marginal distribution shapes depend only weakly on variations of the cuts, see Figs. S4 in SM [70]. The marginal probability density is strongly enhanced for opening angles in the range –, and forms a broad maximum at . With the same range, the dihedral-angle distribution exhibits a peak around – (see Fig. S5 for details). This pattern is suggestive of configurations containing two dineutron-like pairs on opposite sides of the system c.m., with nonzero torsion and dihedral angles indicating twisting and non-planarity. While Ref. [100] suggested symmetric and coplanar “great-circle” configurations in 8He based on a pure shell-model occupation, our results show that such coplanar arrangements are strongly reduced once many-body correlations are treated explicitly. These features suggest that the inter- geometry itself is modified by many-body correlations. Consequently, the four-neutron subsystem favors an extended, mildly twisted configuration that is approximately spatially symmetric about the system c.m. In addition to this dominant extended configuration, tetraneutron-like geometrical components corresponding to the region of small opening angles, where the four neutrons reside on the same side of the system c.m., occur with lower probability. As shown in Fig. 3(b), this component accounts for about 5% of the sampled configurations shown in panel (b) and is concentrated at in the range –. Its associated torsion angle and dihedral angle (see Fig. S5 also lie predominantly in the range –, indicating a strongly noncoplanar and twisted four-neutron geometry compared to the symmetric – configuration. For comparison, the distribution in 8He is shifted to smaller angles and more strongly suppressed at large , indicating a less prominent symmetric - geometry than in 7H under the same selection. This may be attributable to core-induced medium effects on the four-neutron configuration, in that the different proton background and core structure modify the environment experienced by the four neutrons.
@@
To further elucidate these two geometries, Figs. 3(c–f) display the internal angular and radial correlations for configurations corresponding to the extended, approximately symmetric – geometry at and the compact tetraneutron-like geometry at (see Figs. S6 in SM [70] for additional details). For the extended – geometry, Figs. 3(d,f), each dineutron pair retains a small internal opening angle , with its distribution peaked near , and the two pairs exhibit nearly symmetric distributions of the distances (defined as the distances from each pair c.m. to the system c.m.) that extend to about 4–5 fm. In contrast, the tetraneutron-like geometry, Figs. 3(c,e), shows enhanced probability density near characteristic radial-distance pairs fm and fm from the system c.m., indicating two radially asymmetric pairs located on the same side of the system c.m. Their internal opening-angle distributions are shifted to larger values, with a pronounced maximum near and lying predominantly in the range -. These configurations are also accompanied by large torsion and dihedral angles between and (see Fig. S5). Together, these features lead to a non-coplanar, interlaced geometry. This geometry represents a compact, strongly correlated four-neutron configuration, in which the distinction between the two clusters becomes less pronounced, suggesting enhanced four-body correlations among the four neutrons. It may be viewed as a structural precursor to a possible tetraneutron configuration. Overall, these results highlight two representative geometries of the correlated four-neutron subsystem in 7H: an approximately spatially symmetric, extended - configuration with large inter-pair separation and finite torsion, and a compact, twisted tetraneutron-like configuration characterized by enhanced four-body correlations. The extended - geometry dominates the marginal distribution at the level, while the compact component contributes the remaining , with the relative weights only weakly sensitive to reasonable variations of the selection (see Figs. S4). The strong dominance of the extended – geometry in the neutron-rich light nuclei studied here implies that experimental searches for tetraneutron-like configurations are likely to receive sizable contributions from symmetric – modes, complicating the isolation of a genuinely compact, tetraneutron-like signal.
@@
Summary: We perform ab initio NLEFT calculations of the neutron-rich nuclei 6,7H and 8He, including full -body correlations and quantifying uncertainties with a Bayesian analysis over an ensemble of chiral interactions. We predict the ground-state energies of the extreme hydrogen isotopes 6,7H and obtain the energies of the benchmark nucleus 8He consistent with available data, while simultaneously uncovering their internal multi-neutron structure. Intrinsic densities and two-neutron correlation functions indicate a compact triton (alpha) cluster in 7H (8He) surrounded by four valence neutrons that form compact dineutron pairs. Building on this, we evaluate four-body correlations directly on lattice Monte Carlo configurations and construct a reduced four-body correlation function to probe the internal geometry of the four-neutron subsystem. We extract the corresponding spatial and angular correlation patterns between each nucleon in two representative geometries. We find that the 7H ground state is dominated (about 95%) by an approximately spatially symmetric, extended – geometry with large inter-pair separation and finite torsion, while a compact, twisted tetraneutron-like configuration with pronounced four-body correlations accounts for the remaining 5%. These results provide a microscopic picture of emergent multi-neutron correlations in 7H and 8He and suggest that experimental searches for tetraneutron-like signatures may receive non-negligible contributions from symmetric - modes, which can complicate the isolation of a genuinely compact signal.